OpenStudy (anonymous):

PLEASEEE HELPP ??? Tony is trying to calculate how much ice cream he will need to order for his ice cream shop by figuring out the volume of ice cream in one scoop. If the typical scoop of ice cream is 2 inches in radius, what is the volume of one scoop of ice cream? Round to the nearest tenth.

OpenStudy (anonymous):

formula is \[\frac{ 4 }{ 3 }\pi r ^{3}\]

OpenStudy (anonymous):

OKAY, where do i plug in 2 ?

OpenStudy (anonymous):

at r

OpenStudy (anonymous):

since r is equal to the radius

OpenStudy (anonymous):

so its this 4/3(3.14)2^3 ??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what is your answer?

OpenStudy (anonymous):

uhmm Idk what to do first paranthesis ??

OpenStudy (anonymous):

do you already know how to simplify \[2^{3}\]

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you need to multiply 8 with 3.14

OpenStudy (anonymous):

get the product and multiply the product with 4/3

OpenStudy (anonymous):

25.12

OpenStudy (anonymous):

multiply it with 4/3

OpenStudy (anonymous):

what do you get???

OpenStudy (anonymous):

an easy way to multiply fractions with whole numbers is first to divide the whole number with the denominator in this case 3 and multiply the quotient with the numerator in this case 4

OpenStudy (anonymous):

you get 33.493333333.....

OpenStudy (anonymous):

round it to the nearest tenth

OpenStudy (anonymous):

thats the final answer

OpenStudy (anonymous):

wait how did u get that decimal ?

OpenStudy (anonymous):

and is that rounded to the nearest tenth

OpenStudy (anonymous):

wait nvm i see how you git it! but is this answer rounded to the nearest tenth

OpenStudy (anonymous):

that's not the rounded formm,,

OpenStudy (anonymous):

isn't it 30

OpenStudy (anonymous):

can you round it??

OpenStudy (anonymous):

i did i got 30

OpenStudy (anonymous):

25.12 /3= 8.3733

OpenStudy (anonymous):

multiplied by 4

OpenStudy (anonymous):

you get 33.493

OpenStudy (anonymous):

try to use a calculator

OpenStudy (anonymous):

nooo look.

OpenStudy (anonymous):

type the number if its not showing up

OpenStudy (anonymous):

yup,,, 33.5 is the final answer,, 33.49333 is not yet the final answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you must have clicked tens not tenth, tenth is different from tens

OpenStudy (anonymous):

I did click it.

OpenStudy (anonymous):

its just not showing when u press the link

OpenStudy (anonymous):

how would u do this one ? Marissa has a snow cone cart that serves about 30 people each hour at the local beach. If the cone shaped snow cone cups have a height of 3 inches and a radius of 1.5 inches, what volume of shaved ice is needed each hour? Round to the nearest cubic inch.